Math 181  Midterm 1  September 25, 1992

NAME: (PRINT) SIGNATURE: SSN:

The use of calculators is not permitted on this exam. Put all your work in the space provided. NO WORK – – NO CREDIT. Compute as many of the following integrals as time allows.

1. (15 pts) ∫ \frac{x}{x^{2} − 2x + 2} dx

Math 181  Midterm 1  September 25, 1992

2. (15 pts) ∫ \frac{dx}{(1 + x^{2})^{2}}

3. (10 pts) ∫ log(x^{2} + 1)

4. (15 pts) ∫ \frac{x^{2} + 4x + 2}{x(x + 1)^{2}} dx

5. (15 pts) ∫ x sin^{−1} x dx

6. (15 pts) ∫ \frac{x}{\sqrt{16 − x^{4}}} dx 7. (15 pts) ∫ \frac{dx}{\sqrt{x^{2} − 2x}}